OpenStudy (anonymous):

If g(x) = x2 + 3, find g(4)

OpenStudy (theviper):

16+3=?

OpenStudy (anonymous):

Just replace x with 4. Solve from there.

OpenStudy (jackellyn):

plug in the 4 every time you see an x g(4)= 4^2 + 3 = 16 + 3 = ?

OpenStudy (theviper):

put the value of g(x) in eq & get ur answer:)

OpenStudy (hba):

u need to plug in 4 thats all

OpenStudy (hba):

WELCOME TO OPEN STUDY

OpenStudy (anonymous):

the answer is 11

OpenStudy (anonymous):

(4)2 + 3=? 8+3=11

OpenStudy (hba):

@Ashley99 i think its a square not a 2 with the x

OpenStudy (anonymous):

on the side... how do you "plug in" a number? is it like to a wall socket?

OpenStudy (anonymous):

k

OpenStudy (anonymous):

then your answer would be 19.............

OpenStudy (anonymous):

4^2=16 16+3=19 thnx hba

OpenStudy (hba):

@Ashley99 Ur welcome

OpenStudy (anonymous):

Wow thanks I picked 19 at first but wanted to make sure thanks for confirming :D